Now attempt this exercise on depreciation methods:

On January 2, Year 1, Logan Co. purchased a manufacturing machine for $864,000. The machine has an 8-­year estimated life and a $144,000 estimated salvage value. Logan expects to manufacture 1,800,000 units over the life of the machine. During Year 2, Logan manufactured 300,000 units. Calculate the Year 2 depreciation expense using (1) straight-line depreciation, (2) double-declining balance depreciation, and (3) sum-of-the-years digits
Answer #1
Year 2 depreciation
Straight line $      90,000.00
DDB $   1,62,000.00
sum-of-the-years digits $   1,40,000.00

Working

Double declining Method
A Cost $     8,64,000.00
B Residual Value $     1,44,000.00
C=A - B Depreciable base $     7,20,000.00
D Life [in years] 8
E=C/D Annual SLM depreciation $        90,000.00
F=E/C SLM Rate 12.50%
G=F x 2 DDB Rate 25.00%
Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
1 $      8,64,000.00 25.00% $         2,16,000 $   2,16,000.00 $   6,48,000.00
2 $      6,48,000.00 25.00% $         1,62,000 $   3,78,000.00 $   4,86,000.00
Sum-of-the-Years'-Digits Method
A Cost $     8,64,000.00
B Residual Value $     1,44,000.00
C=A - B Depreciable base $     7,20,000.00
Sum of digits with 5 years (8+7+6+5+4+3+2+1) 36
Year Depreciable base Formula Depreciation expense Accumulated Depreciation Ending Book Value
1 $      7,20,000.00 (([720000x8/36]/) $   1,60,000.00 $   1,60,000.00 $   7,04,000.00
2 $      7,20,000.00 (([720000x7/36]/) $   1,40,000.00 $   3,00,000.00 $   5,64,000.00
Straight line Method
A Cost $     8,64,000.00
B Residual Value $     1,44,000.00
C=A - B Depreciable base $     7,20,000.00
D Life [in years left ]                             8
E=C/D Annual SLM depreciation $        90,000.00